From acd@roe.ac.uk Wed Apr 16 10:54:08 2014 Date: Tue, 15 Apr 2014 21:59:27 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tkerr@jach.hawaii.edu" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes, 11 April 2014 WFAU VDFS Science Archives progress meeting minutes, 11 April 2014 Present: MAR, RPB, RSC, MSH, ACD Apologies: RGM, NJC, ETWS, NCH, KTN, AL DoNM: 10am Friday 25 April 2014 in the Villas meeting room, MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: All to read the Twiki pages: http://apache.roe.ac.uk/twiki/bin/view/WFAU/BackupReview Done. ACTION: ALL to add/maintain entries on http://apache.roe.ac.uk/twiki/bin/view/WFAU/ForthcomingDeadlines Discharged: a review of forthcoming deadlines would be added as a regular agenda item. ACTION: RPB to make GPS public in DR8. Done. ACTION: MAR to supply list of non-surveys that could be released. Done. Actions partly discharged but continuing: ---------------------------------------- None. Actions carried forward from 14/3/2014 meeting: ---------------------------------------------- ACTION: ACD, RPB to formulate a policy for backups and for checking that data could be recovered from the backups. Continues. Notes on a proposed policy are available on the Twiki and discussion is still in progress. One associated item that remained to be looked at was the options for taking a nightly backup of the SVN repository. An alternative approach might be for WFAU to run its own SVN repository. ACTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Continues. ACTION: RSC to investigate modifying CU7 to recover ATLAS VST unmatched sources around frameset boundaries. Continues. Investigations had begun and a strategy had been worked out. However, work is currently suspended pending news from the ATLAS meeting as to whether we should proceed with the suggested approach, given its less than satisfactory consequences and our lack of resources. ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy (http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan) but was not present to report progress. ACTION: NJC to sort out the magnitude limits for V1.3 data. Continues. NJC was not present. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. Specific points and new actions: ------------------------------- Project management: * The one-year extension to the WFAU rolling grant was now in place. WFCAM, VISTA and VST updates: * The transfer of UKIRT to new ownership had been delayed and was still pending. Comments and issues arising from CASU minutes: * No minutes had been received since the previous meeting. Networking: * Nothing to report (ETWS was away). WSA/VSA/OSA/GES Operations: * RPB reported that a VMC release had crashed (towards the end of the process or preparing a release). Investigation was deferred until NJC's return. * ETWS was away and so could not report on progress. Hardware and Systems: * MSH reported: - The new NAS box does not work and the motherboard is to be replaced. - The new Web server is up and running and answers to the name of 'unas'. - RAMESES2 died and was resurrected. (Once again fending off the serpent Apep and completing its voyage in the night-boat through the hours of darkness?) - The changes necessary to fix the heartbleed security loophole had been applied to the Web server. - Windows has been installed on the test machine so that tests can be performed on the borrowed Nytro RAID card. This card will have to be returned next week, so the tests are urgent. - A 10 Gbit ethernet switch and associated cards had arrived, also for testing. Survey Data Release: * RPB noted the following survey releases: VHS P90 v1.3 and VIKING P90. * MAR reported that ESO had finally accepted the VMC release. The VVV release had largely been accepted, but some minor changes were required to VARCAT. Software: * RSC reported that VVV performance testing is in progress, with the focus currently being on the Nytro RAID array and what improvements it can provide; when our loan of it is over effort will revert to investigating some software trickery. Progress updates are on the TWiki: http://apache.roe.ac.uk/twiki/bin/view/WFAU/VVVReleasePerformance Non-survey Data Release: * None. Astrogrid deployment, VO & Data Analysis services: * MSH noted that he had added GPS to DR8. Miscellaneous: * ACD reported that the two new Gaia posts had been sent to Physics and Astronomy HR for submission for advertisement by the University. Note added to minutes: the jobs are now available on the University's Jobs Web site. ACD, original: 11/4/14, revised: 15/4/14. -- Scanned by iCritical.